The supervisor sells 1 type of tea for $11.00 pd and the other for $5.00 pd The supervisor wants to make 70 pds of the $11 tea for $7.00 pd....how much of the $5.00 should be used?
Answer by Theo(13342)

let x = the number of pounds of the 11 dollars a pound tea.
let y = the number of pounds of the 5 dollars a pound tea.
x + y = 70 pounds
11x + 5y = 7*70 = 490 dollars
use the first equation to solve for y to get y = 70 - x
substitute for y in the second equation to get:
11x + 5(70-x) = 490
simplify this equation to get:
11x + 350 - 5x = 490
combine like terms and subtract 350 from both sides of the equation to get:
6x = 140
divide both sides by 6 to get:
x = 23.3333333
if x = 23.33333333 then y = 46.66666667 because x + y = 70
when x = 23.33333333 and y = 46.66666667, the equation of:
11x + 5y = 490 becomes:
11*23.33333333 + 5*46.66666667 = 490 which becomes:
256.66666666 + 233.33333334 = 490.
numbers check out so the answer is:
46 and 2/3 pounds of the 5 dollar a pound tea needs to be mixed with 23 and 1/3 pounds of the 11 dollar tea.